OpenStudy (anonymous):

//fbcdn-sphotos-b-a.akamaihd.net/hphotos-ak-frc3/1377338_544606995617879_1856972893_n.jpg @AllTehMaffs

OpenStudy (anonymous):

hallo

OpenStudy (anonymous):

Can you help me with this problem ?

OpenStudy (anonymous):

V = IR if I put this got 2.5 A

OpenStudy (anonymous):

the equation for current in the circuit with a capacitor is \[I(t) = \frac{Vo}{R}e^{-\frac{t}{\tau_o}}\] So if t is really large, that term should go to zero, and current should just look like \[I(t) = \frac{Vo}{R}\]

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

my pleasure :)